# Scaling Methods To Estimate Macroscopic Fundamental Diagrams in Urban Networks with Sparse Sensor Coverage

### v1

@@ 5.2 Comparison of proposed methods
The results in Figure 6 and Figure 7, combined with the RMSE values presented in Table 2, provide insight into the comparative performance of the proposed statistical (non-hierarchical, 2-hierarchy (2-H), 3-hierarchy (3-H)) and the variogram methods (non-hierarchical) for estimating network flow under different LD coverage percentages (5%, 10%, 20%, and 30%). The temporal flow estimations in Figure 6 demonstrate that higher LD coverage leads to more accurate flow predictions across all methods. This trend is further supported by the RMSE values, which generally decrease as LD coverage increases, especially for the hierarchical and Variogram methods.
At the lowest LD coverage (5%), the Variogram method is unable to estimate flow across the entire network, reflecting a limitation in this approach when LD coverage falls below 10%. Meanwhile, the remaining methods, particularly the 3-Hierarchy, continue to provide estimations, though with elevated RMSE values (e.g., RMSE of 65-64 on day 1 and day 4 for the 3-Hierarchy method). This highlights the robustness of hierarchical frameworks compared to geospatial estimation techniques like the Variogram, which require a minimum density of LDs for effective network-wide prediction. At 10% LD coverage, the Variogram method becomes viable and achieves relatively low RMSE values compared to non-hierarchical methods, indicating an improvement in estimation accuracy with increased LD density. Across all methods, higher LD coverage continues to yield more accurate predictions, with RMSE values substantially decreasing by the time LD coverage reaches 30%. Here, both the 3-Hierarchy and Variogram methods achieve RMSE values close to or below 15, showcasing their potential for precise flow estimation even with partial network monitoring. In the statistical methods, increasing the number of hierarchical levels in the network consistently enhances estimation accuracy. For any given LD coverage percentage, the 3-hierarchy (3-H) network consistently outperforms both the 2-hierarchy (2-H) and non-hierarchical approaches.
@@
These results underscore the robustness of hierarchical models, particularly in sparse LD conditions where unstructured methods like the Variogram may falter. The 3-Hierarchy method, in particular, shows resilience across all coverage levels, while the Variogram method performs strongly at 10% and higher LD coverage, demonstrating its potential for accurate network-wide estimations when sufficient spatial data is available. The non-hierarchical approach, while less accurate, serves as a baseline, illustrating the limitations of unstructured models in low-coverage scenarios. Overall, these findings underscore the value of incorporating hierarchical and spatially structured models to enhance traffic flow estimations in partially monitored networks.
@@ 5.3 Comparing MFDs
This section demonstrates network MFDs estimated from the partially equipped network using the proposed scaling methodologies and compares them with the actual MFDs from the fully equipped network. In section 3.1, we proposed the network flow $(\hat{q})$ and density $(\hat{k})$ using the statistical scaling method, and section 3.2 described $\hat{q}$ estimation using the variogram. Similarly, we can estimate $\hat{k}$ using variogram methods by estimating detector-level density.
The estimated macroscopic fundamental diagrams (MFDs) derived from the proposed methodologies across varying LD coverage percentages demonstrate clear differences in accuracy among the methods employed, as summarized in Figure 8 and Table 3. The 5% LD scenario represents the most challenging condition for flow estimation due to sparse detector data. Among the methods, the three-hierarchy (3-H) statistical scaling achieves the lowest root mean square error (RMSE) of 48.9 veh/h, significantly outperforming the two-hierarchy (2-H) scaling (51.8 veh/h) and the non-hierarchical approach (175.5 veh/h). The high coefficient of determination ($R^{2}=0.97$) for the 3-H method further highlights its robustness in low-data scenarios, maintaining strong alignment with the actual MFD. For 10% LD, the 3-H scaling method again delivers superior results with an RMSE of 45.3 veh/h, slightly better than the 2-H method (45.7 veh/h), variogram (55.2 veh/h), and markedly better than the non-hierarchical approach (137.6 veh/h). At 20% LD, the differences between the hierarchical methods narrow further, with 3-H achieving 36.5 veh/h RMSE compared to 37.3 veh/h for 2-H. Both methods significantly outperform the non-hierarchical method (133.5 veh/h), demonstrating the effectiveness of incorporating spatial structure in flow estimation. At the highest detector coverage of 30%, all methods perform well due to the availability of more extensive data. However, the 3-H method remains the most accurate, with an RMSE of 35.9 veh/h, compared to 37.1 veh/h for 2-H and 39.9 veh/h for the non-hierarchical method. The variogram method also performs well in this scenario, achieving an RMSE of 51.5 veh/h, though its application is limited to LD coverage of 10% or greater.
@@
6 Discussion
The results underscore the critical role of hierarchical scaling in improving flow and MFD estimation accuracy, particularly under sparse data conditions. The 3-H method consistently outperforms the 2-H and non-hierarchical approaches across all LD levels, with its advantage most pronounced at low LD coverage (5% and 10%). This is likely due to the additional spatial structure incorporated in the 3-H scaling, which allows it to better capture localized variations in flow and density. At higher LD coverage (20% and 30%), the benefits of hierarchical approaches persist, though the performance gap between methods narrows. The non-hierarchical method shows improved accuracy as more data becomes available, but it remains less reliable than its hierarchical counterparts due to its inability to effectively extrapolate flow patterns in under-monitored regions. Interestingly, the variogram approach demonstrates competitive performance for LD coverage levels of 10% or more, but it cannot be applied at 5% LD due to insufficient data for spatial interpolation. This limitation highlights the trade-offs between different methods in terms of data requirements and estimation capabilities. Overall, the findings emphasize the importance of leveraging hierarchical frameworks and structured statistical methods for MFD estimation, particularly in scenarios where data availability is limited. The superior performance of the 3-H method highlights its potential for practical applications in network flow analysis, even under challenging conditions. The insights gained from this study can inform future efforts to optimize traffic monitoring systems and enhance flow prediction accuracy.
@@ 7 Conclusion
This study presents a robust methodology for estimating network-wide traffic flow characteristics in partially sensor-equipped urban networks. By integrating statistical scaling and geospatial imputation techniques, the framework provides accurate estimations of macroscopic fundamental diagrams (MFDs), even in networks with sparse detector coverage. The hierarchical network approach, which categorizes links into three levels based on spatial and functional attributes, consistently outperforms non-hierarchical approaches, particularly at low levels of detector deployment (5%–10% LD). The results demonstrate that three-hierarchy statistical scaling achieves the lowest RMSE and highest $R^{2}$ values, highlighting its reliability for sparse sensor networks.
While the variogram-based geospatial method captures spatial correlations effectively, its limitations are evident in networks with very low detector density $(<10\%LD)$. As a geospatial interpolation tool, the variogram model estimates non-equipped link variables based on spatial distribution. Although route distances between links were used instead of Euclidean spatial distances, some interpolation errors may arise, affecting accuracy. In contrast, the statistical method is more promising because it assumes that links within the same hierarchy exhibit similar mean and covariance properties. However, challenges may emerge due to deviations in link characteristics, which can undermine the consistency of the scaling factors across different scenarios.
Despite these limitations, the proposed methodology offers a scalable and practical solution for estimating network-wide traffic conditions under varying levels of sensor availability. The findings underscore the potential of incorporating network structure and spatial dependencies in traffic modeling for real-world applications. This study bridges the gap between partial sensor data and full-network traffic estimation, providing a scalable solution for cities with limited sensor coverage.

### v2

@@ 4.2 Comparison of proposed methods
The results presented in Figures 6 and 7, along with the RMSE values in Table 2, provide a detailed comparison of the proposed hierarchical network scaling method, the variogram-based approach, and the baseline non-hierarchical network scaling. This evaluation focuses on their performance in estimating network flow under different levels of LD coverage, specifically at 5%, 10%, 20%, and 30%. The temporal flow estimations in Figure 6 show that the baseline (non-hierarchical) quickly loses accuracy when the LD coverage is reduced below 20%. It results in an inaccurate estimation of network flow, causing significant bias in the MFD. Recall the usual network with 5.78 % LDs coverage, which is very low and far from 20% LDs coverage. At the lowest LD coverage (5%), the Variogram method is unable to estimate flow across the entire network, reflecting a limitation in this approach when LD coverage falls below 10%. At 10% LD coverage, the Variogram method becomes viable and achieves relatively low RMSE values compared to non-hierarchical network scaling. Although the variogram method with 10% and 20% LD coverage demonstrates improved accuracy compared to the baseline, it also results in greater fluctuations in higher flow regimes (see Figure 6). Meanwhile, the hierarchical network scaling performs better in the low coverage of LD with minimum deviation and fluctuation compared to the actual flow. In the hierarchical method, the 3-H network cluster demonstrates best among the overall methodologies. The 3-Hierarchy continues to provide estimations, though with elevated RMSE values (e.g., RMSE of 65-64 on day 1 and day 4 for the 3-Hierarchy method). This highlights the robustness of hierarchical frameworks compared to geospatial imputation techniques like the Variogram, which require a minimum density of LDs for effective network-wide prediction compared to the non-hierarchical scaling approach. Across all methods, higher LD coverage consistently leads to more accurate predictions. RMSE values significantly decrease once LD coverage reaches 20%-30%, a benchmark that is challenging to achieve in practice. The average RMSE for five days of network traffic flow estimation at LD coverage of 10% is 29.2 vehicles per hour (veh/h) using the 3-H approach, compared to 41.6 veh/h with the variogram method. When the LD coverage decreases to 5%, the RMSE for the 3-H method increases to 56 veh/h, which remains a reasonable estimation given the reduced coverage of LDs. In the hierarchical network methods, increasing the number of hierarchy groups in the network consistently enhances estimation accuracy. Therefore, at any given LD coverage percentage, the 3-hierarchy (3-H) network consistently outperforms the baseline and other methods.
@@
These results underscore the robustness of hierarchical network scaling, particularly in sparse LD conditions where unstructured methods like the Variogram may falter. The 3-H method, in particular, shows resilience across all coverage levels, while the Variogram method performs strongly at 10% and higher LD coverage, demonstrating its potential for accurate network-wide estimations when sufficient spatial data is available. The baseline approach, applied under the lower usual case of 5% LD coverage, yields a network-average flow estimation with an RMSE of 234 veh/h. This value is notably four times higher than the error achieved using the proposed three-level hierarchical (3-H) network scaling method. These findings underscore the importance of adopting a hierarchical network approach to improve the accuracy of traffic variable estimation at the network level.
@@ 4.3 Comparing MFDs
This section demonstrates network MFDs estimated from the partially equipped network using the proposed scaling methodologies and compares them with the actual MFDs from the fully equipped network. In section 2.1, we proposed the network flow $(\hat{q})$ and density $(\hat{k})$ using the hierarchical scaling method, and section 2.2 described $\hat{q}$ estimation using the variogram. Similarly, we can estimate $\hat{k}$ using variogram methods by estimating detector-level density.
The estimated MFDs derived from the proposed methodologies across varying LD coverage percentages demonstrate clear differences in accuracy among the methods employed, as summarized in Figure 8 and Table 3. The 5% LD scenario represents the most challenging condition for MFD estimation due to sparse detector data. Among the methods, the three-hierarchy (3-H) statistical scaling achieves the lowest RMSE of 48.9 veh/h, significantly outperforming the two-hierarchy (2-H) scaling (51.8 veh/h) and the non-hierarchical approach (175.5 veh/h). The high coefficient of determination ($R^{2}=0.97$) for the 3-H method further highlights its robustness in low-data scenarios, maintaining strong alignment with the actual MFD. For 10% LD, the 3-H scaling method again delivers superior results with an RMSE of 45.3 veh/h, slightly better than the 2-H method (45.7 veh/h), variogram (55.2 veh/h), and markedly better than the non-hierarchical approach (137.6 veh/h). At 20% LD, the differences between the hierarchical methods narrow further, with 3-H achieving 36.5 veh/h RMSE compared to 37.3 veh/h for 2-H. Both methods significantly outperform the non-hierarchical method (133.5 veh/h), demonstrating the effectiveness of incorporating spatial structure in flow estimation. At the highest detector coverage of 30%, all methods perform well due to the availability of more extensive data.
@@
Using the hierarchical network clustering methods, we estimate the full network MFD, as illustrated in Figure 9, and compare these results with the baseline method. The full network consists of all 2456 links as described in data description section. It is evident from the figure that the baseline method underestimates the MFD, particularly in the critical density regime. However, in the free-flow regime, both approaches yield comparable results. Recalling the network LD coverage as 5.68%, we can refer to the MFDs in Figure 8 (a) for 5% LD; the same pattern is observed here. To simplify our analysis, we fitted the scattered MFDs using a non-linear model and created 95% confidence intervals. The best-fit three-hierarchy model can estimate about 66 veh/h more in the critical density regime, along with a slight improvement in the maximum observed network density. Although it is not too significant for our case, for a larger and more hierarchical network, it may significantly impact the MFD estimation. While the limited LD coverage poses challenges for validating the full network MFD, the hierarchical scaling approach demonstrates a significant improvement over the baseline method, particularly in capturing the critical density region.
Figure 9: Comparing the estimated MFD using a hierarchical network cluster with three hierarchies to the baseline non-hierarchical method for the entire network consisting of 2456 links.
@@ 5 Conclusion
This study presents a novel methodology for estimating network-wide traffic characteristics in partially sensor-equipped urban networks. By integrating hierarchical network scaling and geospatial imputation techniques, the framework provides accurate estimations of MFDs, even in networks with sparse detector coverage. The results underscore the critical role of hierarchical scaling in improving network traffic variables and MFD estimation accuracy, particularly under sparse data conditions. The 3-H method consistently outperforms (lowest RMSE and highest $R^{2}$ values) the 2-H and variogram approaches across all LD levels, with its advantage most pronounced at low LD coverage (5% and 10%). This is likely due to the additional spatial structure incorporated in the hierarchical scaling, which allows it to better capture localized variations in flow and density. Also, we found that more number of network hierarchies influence better estimation of network traffic states and MFDs. The estimation of the full network’s MFD using three hierarchical network clusters reveals significant deviations in network saturation flow and critical density range compared to traditional methods, as illustrated in Figure 9. The baseline scaling method shows improved accuracy as more data becomes available, but it remains less reliable than its hierarchical counterparts due to its inability to effectively extrapolate flow patterns in under-monitored regions. Interestingly, the variogram approach demonstrates competitive performance for LD coverage levels of 10% or more with small fluctuations in high flow regimes, but it was incompetent at lower (5%) LD due to insufficient data for spatial imputation. This limitation highlights the trade-offs between different methods in terms of data requirements and estimation capabilities. As a geospatial interpolation tool, the variogram model estimates non-equipped link variables based on spatial distribution. Although route distances between links were used instead of Euclidean spatial distances, some interpolation errors may arise, affecting accuracy. In contrast, the hierarchical method is more promising because it assumes that links within the same hierarchy exhibit similar mean and covariance properties. It offers a scalable and practical solution for estimating network-wide traffic conditions under varying levels of sensor availability. The findings underscore the potential of incorporating network structure and spatial dependencies in traffic modeling for real-world applications. Therefore, this study bridges the gap between partial sensor data and full-network traffic estimation for cities with limited sensor coverage. Overall, the findings emphasize the importance of leveraging network hierarchical frameworks for MFD estimation, particularly in scenarios where data availability is limited. The superior performance of the 3-H method highlights its potential for practical applications in network flow analysis, even under challenging conditions.